ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: 53 Kashiwada, Toyoko, Saitama College, Japan The integrability of almost generalized Hopf manifolds An almost generalized Hopf structure is a locally conformal Kähler structure whose Lee form is parallel. An integrability condition of almost generalized structures will be presented, and be applied to some classes of almost generalized Hopf structures.